Roko paid $12.50 for 25 game tickets.Louisa paid $17.50 for 35 game tickets.What is the constant of proportionality?


1
Expert's answer
2021-08-26T08:36:45-0400

If Roko paid $12.50 for 25 game tickets anf Louisa paid $17.50 for 35 game tickets, then the constant of proportionality is:

12.5/25 = 17.5/35 = $0.5 per ticket.